# The Inadequacy of Similarity-based Privacy Metrics: Privacy Attacks against “Truly Anonymous” Synthetic Datasets

### v2

@@ 6 Experimental Evaluation
In this section, we evaluate the effectiveness of our privacy attacks. First, we show DifferenceAttack’s perfect performance in membership and attribute inference and compare it to existing methods. We then demonstrate that ReconSyn successfully recovers the train outliers in different settings. More specifically, we measure the performance of ReconSyn (SampleAttack and SearchAttack) against PrivBayes, MST, DPGAN, PATE-GAN, and CTGAN (all in a non-private setting, i.e., $\epsilon=\infty$) on increasingly more complex datasets (2d Gauss, Adult Small, Adult, Census, MNIST, described in Appendix B).
@@ 6.1 DifferenceAttack
We start by demonstrating that the simple DifferenceAttack achieves better results compared to state-of-the-art attacks (sota) with considerably less computation (see Table 2) owing to the leakage coming from the privacy metrics.
@@ 6.1.1 Membership Inference
To estimate the AUC from GroundHog [115], Querybased [63], and DOMIAS [132], we use 1,200 models (for Groundhog and Querybased) to train the classifier and 800 models (for all three attacks) for testing when attacking PrivBayes and MST, and 300 for training/200 for testing against the GANs, as they take much longer to converge. This is comparable to Groundhog’s default 1,000 and Querybased’s 3,500 models, even though we test more numerous and more complex generative model implementations. (While we will include AUC results with the default number of models, we do not expect scores to significantly change as they mostly plateau.) As shown in the first three columns in Table 2, with a couple of exceptions (Groundhog vs. DPGAN and Querybased vs. DPGAN and CTGAN), the three sota attacks achieve AUC results below 0.66 and can take over 120 hours to run, particularly against the GANs.
@@ 6.1.2 Attribute Inference
By contrast, sota attribute inference attacks do not achieve very high AUC scores, being reported as at most 0.75 on PrivBayes [6] and 0.5 for MST [63] and CTGAN [63, 6] ($\epsilon=\infty$). Thus, we do not see the benefit of reproducing experiments with the sota attribute inference attacks. Similarly, we do not adapt the three sota membership inference attacks to attribute inference via multi-membership inference would require hundreds of hours and is unlikely to yield stronger results anyway.
@@ 6.2.1 SampleAttack with One Call
Figure 6: Proportion of reconstructed train outliers for increasing rounds by SampleAttack, Adult Small, Adult, and Census.
@@ 6.2.2 SampleAttack with Multiple Calls
2d Gauss. For 2d Gauss, we attack the “oracle” (introduced in CE3 in Sec. 4), which possesses knowledge of the generative process. Even though no generative model has been exposed to train data, and the oracle has no memory of the synthetic data it generated, SampleAttack reconstructs 95% of train outliers due to leakage from the privacy metrics.
@@
Figure 7: CDF of similarity distances between train and test/synthetic outliers by SampleAttack, MNIST.
@@ 8.1 Summary
In a way, one could compare SBPMs to the inadequate privacy guarantees offered by Diffix-like anonymization systems [106, 42, 20]. Although the functionalities are different (Diffix provides answers to queries), they both allow for a potentially unbounded number of queries while not implementing robust privacy mechanisms like DP. ultimately leading to severe privacy violations.
@@
Competing Interests. Georgi Ganev is a research scientist at Hazy, a provider of synthetic data technologies. Emiliano De Cristofaro is an academic researcher with no competing interests in this work.
@@ 8.2.5 Preventing Overfitting/Memorization
Privacy attacks against synthetic data are generally thought to be caused by generative models memorizing and/or overfitting to the train data [15, 135, 57]. However, we argue that reconstruction attacks would still be possible in the context of SBPM-based techniques even if one could somehow guarantee that the underlying models do not overfit or memorize the data.
@@ 8.3 Further Considerations
Remarks on DP. With end-to-end DP pipelines, privacy becomes an attribute of the process [131], and, as per the post-processing property, any synthetic data sample is also DP. DP also provably bounds the risks of singling out predicates [21] and of linkability and inference attacks [47]. However, our review of the product offerings shows that DP is not the standard; companies in this space rarely use it and, in some cases, seem to argue against it [83, 98].
@@
Limitations and Future Work. ReconSyn successfully reconstructs most outliers in various settings, yet it could still be optimized. In future work, we plan to investigate the minimum number of records the adversary needs to generate as well as the effect of preventing the adversary from augmenting the generated synthetic data. Follow-up research could also study the impact of privacy metrics using continuous distances, which could make the calculations more precise and yield better search algorithms, or extend the evaluation to additional synthetic data techniques based on Diffusion Models or Transformer architectures.
@@ Appendix A Reconstruction of All Train Data
Table 4: Reconstruction of any train records by ReconSyn, Adult.
@@ Appendix B Datasets
We describe the datasets used for our evaluation and our criteria for defining outliers; see Table 5. We construct two controllable datasets (2d Gauss and 25d Gauss) based on the normal distribution and use three “standard” datasets after some sampling/pre-processing (Adult, Census, and MNIST).
@@
Outliers Definition. While various definitions of outliers exist in literature [13, 79], we define underrepresented data regions and outliers in a way that captures various scenarios, aiming for an intuitive selection of roughly 10% of the train data (to serve as our targets, as noted in Table 5). For the 2d Gauss dataset, we identify outliers as points lying beyond a certain distance from the center. In the Adult Small, Adult, and Census datasets, outliers are the smallest clusters determined by a Gaussian Mixture model. For MNIST, the digit 0 is deliberately underrepresented. We apply the same strategy or fitted model to the test data. UMAP’s distance-preserving feature allows these outlier identification strategies to be visually verified in Fig. 10(a)10(d).
@@ Appendix D Additional SBPMs Counter-Examples
Alternatively, if we fix a synthetic dataset, randomly split the available sensitive data into 50/50% train/test datasets, and feed them into the metrics, we again run into inconsistencies. Out of 1,000 repetitions, only 380 instances pass all three tests. This highlights the inherent randomness in the train/test split, which incorporates instability into the evaluation process.

### v3

@@ 6 Experimental Evaluation
In this section, we evaluate the effectiveness of our privacy attacks. First, we show DifferenceAttack’s perfect performance in membership and attribute inference and compare it to existing methods. We then demonstrate that ReconSyn successfully recovers the train outliers in different settings. More specifically, we measure the performance of ReconSyn (SampleAttack and SearchAttack) against PrivBayes, MST, DPGAN, PATE-GAN, and CTGAN (in non-private settings, i.e., $\epsilon=\infty$) on increasingly more complex datasets (2d Gauss, Adult Small, Adult, Census, MNIST, described in Appendix B).
@@ 6.1 DifferenceAttack
We start by demonstrating that the simple DifferenceAttack achieves better results compared to state-of-the-art attacks (SOTA) with considerably less computation (see Table 2) owing to the leakage coming from the privacy metrics (and their stronger implicit threat model).
@@ 6.1.1 Membership Inference
To estimate the AUC from GroundHog [110], Querybased [61], and DOMIAS [127], we use 1,200 models (for Groundhog and Querybased) to train the classifier and 800 models (for all three attacks) for testing when attacking PrivBayes and MST, and 300 for training/200 for testing against the GANs, as they take much longer to converge. This is comparable to Groundhog’s default 1,000 and Querybased’s 3,500 models, even though we test more numerous and more complex generative model implementations. (We do not expect the AUC scores to significantly change if we used the default number of models, as results mostly plateau.) As shown in the first three columns in Table 2, with a couple of exceptions (Groundhog vs. DPGAN and Querybased vs. DPGAN and CTGAN), the three attacks achieve AUC results below 0.66 and can take over 120 hours to run, particularly against the GANs.
@@ 6.1.2 Attribute Inference
By contrast, SOTA attribute inference attacks do not achieve very high AUC scores, being reported as at most 0.75 on PrivBayes [6] and 0.5 for MST [61] and CTGAN [61, 6] ($\epsilon=\infty$ for PrivBayes/MST). Thus, we do not see the benefit of reproducing experiments with the SOTA attribute inference attacks. Similarly, we do not adapt the three membership inference attacks to attribute inference via multi-membership inference would require hundreds of hours and is unlikely to yield stronger results anyway.
@@ 6.2.2 SampleAttack with Multiple Calls
2d Gauss. For 2d Gauss, we attack the ‘oracle’ (introduced in CE3 in Sec. 4), which possesses knowledge of the generative process. Even though no generative model has been exposed to train data, and the oracle has no memory of the synthetic data it generated, SampleAttack reconstructs 95% of train outliers due to leakage from the privacy metrics.
@@
Figure 6: Reconstruction success rate of train outliers for increasing SampleAttack rounds, Adult Small, Adult, and Census.
Figure 7: CDF of similarity distances between outliers in train-test and train-synthetic datasets observed by SampleAttack, MNIST.
@@ 8.1 Summary
In a way, one could compare SBPMs to the inadequate privacy guarantees offered by Diffix-like anonymization systems [101, 40, 19]. Although the functionalities are different (Diffix provides answers to queries), they both allow for a potentially unbounded number of queries while not implementing robust privacy mechanisms like DP ultimately leading to severe privacy violations.
@@
Competing Interests. This work was done while GG was with Hazy, a synthetic data provider. EDC is an academic researcher with no competing interests in this work.
@@ 8.2.5 Preventing Overfitting/Memorization
Privacy attacks against synthetic data are generally correlated with generative models memorizing and/or overfitting  [15, 130, 55]. However, we argue that reconstruction attacks would still be possible in the context of SBPM-based techniques even if one could somehow guarantee that the underlying models do not overfit or memorize the data.
@@ 8.3 Further Considerations
Remarks on DP. With end-to-end DP pipelines, privacy becomes an attribute of the process [126], and, as per the post-processing property, any synthetic data sample is also DP. DP also provably bounds the risks of singling out predicates [20] and of linkability and inference attacks [45].
@@
Limitations and Future Work. ReconSyn successfully reconstructs most outliers in various settings, yet it could still be optimized. In future work, we plan to investigate the minimum number of records the adversary needs to generate as well as the effect of preventing the adversary from augmenting the generated synthetic data. Follow-up research could also study the impact of privacy metrics using continuous distances, which could make the calculations more precise and yield better search algorithms.
Acknowledgments. We are grateful to the IEEE S&P shepherd and reviewers for their valuable feedback, which helped us to improve our paper. We also thank Theresa Stadler and Yves-Alexandre de Montjoye for their helpful comments.
@@ Appendix A Reconstruction of All Train Data
Table 4: Reconstruction rate of any records by ReconSyn, Adult.
@@ Appendix B Datasets
We describe the datasets used for our evaluation and our criteria for defining outliers; see Table 5. We construct two controllable datasets (2d Gauss and 25d Gauss) based on the normal distribution and use three ‘standard’ datasets after some sampling/pre-processing (Adult, Census, and MNIST).
@@
Outliers Definition. While various definitions of outliers exist in literature [13, 77], we define underrepresented data regions and outliers to capture various scenarios, aiming to intuitively select roughly 10% of the train data (to serve as targets, as noted in Table 5). For 2d Gauss, we identify outliers as points lying beyond a certain distance from the center. In Adult Small, Adult, and Census, outliers are the smallest clusters determined by a Gaussian Mixture model. For MNIST, the digit 0 is deliberately underrepresented. We apply the same strategy/model to the test data. UMAP’s distance-preserving feature allows these outlier identification strategies to be visually verified in Fig. 10(a)10(d).
@@ Appendix C Additional Fundamental Issues of SBPMs
#Cols #Outliers #Synth #Bins #Bins Attack
Datasets Outliers Outliers Success
(* $10^{3}$) Total Empty
2 108 8 442 108 1.00
3 118 4,059 12,632 140 0.84
4 110 3,520 340,475 26,236 0.01
5 95 757 8,915,441 4,573,841 2.08*$10^{-5}$
Table 6: Similarity Filter and reconstruction success rate (12 hour limit), Gauss with different dimensions.
@@ Appendix D Additional SBPMs Counter-Examples
CE3. SF & Reconstruction (continued). We expand CE3 by extending Gauss to higher dimensions and present the results in Table 6. While we successfully isolate all outliers for two dimensions and 84% for three, the number of bins potentially containing outliers grows exponentially, reaching nearly 350,000 for four dimensions and 9,000,000 for five. Consequently, the attack slows down significantly; within 12 hours of runtime, we can only ‘visit’ a small fraction of these bins, rendering the attack ineffective.
Conversely, targeting a single data record, as in standard membership inference attacks, makes the attack more effective. Suppose the target lies at the origin (the zero vector) with no training records nearby. Without conditional data generation, i.e., we naively sample from an oracle, we can create a ‘hole’ around the target for up to 8 dimensions in less than 3 hours. Beyond this, the neighborhood’s dimensionality grows too large. However, with conditional data generation, we can construct the target’s neighborhood one column at a time, conditioning on zeros for the remaining columns. This reduces complexity from exponential to linear, enabling us to encircle the target even in dimensions beyond 100 within seconds.
@@
Alternatively, fixing a synthetic dataset and randomly splitting the sensitive data 50/50 into train/test sets still leads to inconsistencies. Out of 1,000 repetitions, only 380 instances pass all three tests. This highlights the inherent randomness in the train/test split, which incorporates instability into the evaluation process.